# Magic Gems: A Polyhedral Framework for Magic Squares Connecting Combinatorics, Geometry, and Linear Algebra

### v1

@@ 5.1.1 The Vector Interpretation
This interpretation extends to the diagonal conditions. The diagonal covariances $\mathrm{Cov}(D_{\mathrm{main}},Z)$ and $\mathrm{Cov}(D_{\mathrm{anti}},Z)$ measure whether diagonal positions predict value. For magic squares, all four directional indicators—horizontal position, vertical position, main diagonal membership, and anti-diagonal membership—are uncorrelated with value. The magic energy $E(S)$ is simply the sum of squared correlations with these four directions.
@@ 5.1.2 Physical and Algebraic Interpretations
From an algebraic perspective, the zero-covariance condition is equivalent to the defining property of magic squares—equal row and column sums—expressed in the language of moments. The first moment (mean) of each coordinate vanishes by construction; the zero-covariance condition asserts that certain second moments (cross-terms) also vanish. This suggests a hierarchy: might there be higher-order “super-magic” squares characterized by vanishing higher moments?
@@ 5.2 Magic Squares as Ground States
The energy landscape perspective developed in Section 3.7 offers perhaps the most illuminating interpretation of our results. The covariance energy $E(S)$ provides a natural “imbalance” measure for arrangements, and magic squares emerge as the unique configurations where this imbalance vanishes completely.
@@
The dynamical systems interpretation is equally rich. Each magic square defines a basin of attraction in the discrete configuration space, though these basins have measure zero. The comprehensive perturbation analysis (Section 2) confirms that there are no “flat directions” around magic configurations: across 107,388 tested perturbations of 923 magic squares, every single swap increases the energy, supporting the conclusion that magic squares are strict local minima.
The perturbation analysis reveals additional structure within the magic square population. For $n=3$, all eight magic squares exhibit exactly the same minimum perturbation gap ($\Delta_{3}=0.0988$), a consequence of the $D_{4}$ symmetry preserving the embedding geometry. More surprisingly, the 880 $n=4$ magic squares cluster into three distinct bands of perturbation resistance, suggesting the existence of geometric subclasses with different stability properties—a finding that may correlate with convex hull configurations and merits further investigation. The minimum gap scales approximately as $\Delta\propto 1/n^{2}$, quantifying how the energy wells become progressively shallower as order increases.
@@ 5.3 Scaling Laws and Landscape Evolution
Despite the dramatic growth in configuration space ($n!^{2}$ arrangements), the landscape entropy remains approximately constant ($\approx 4.4$$5.0$). This invariance suggests that the relative difficulty of local optimization is independent of scale: the ruggedness of the landscape, measured by the distribution’s shape rather than its position, does not fundamentally change with $n$. This has practical implications for algorithm design: local search heuristics that work for small orders may transfer to larger ones, with the primary challenge being the absolute energy scale rather than landscape structure.
Isolation Persistence.
Across over 460 million sampled arrangements—including exhaustive enumeration for $n=3$, 60 million samples for $n=4$, and 400 million for n=5n=5—no false-positive zero-energy configurations were discovered. While this does not constitute a proof, it provides compelling computational evidence that the Energy Characterization Theorem (Section 3.13) extends beyond the analytically tractable cases. The probability of discovering a magic square by random sampling decreases super-exponentially with $n$, from roughly $1$ in 45,00045{,}000 for $n=3$ to less than $1$ in $10^{17}$ for $n=5$. This quantifies the “needle in a haystack” intuition with concrete numbers.
@@ 5.4 Even versus Odd Orders
Despite these differences, magic squares of all orders exhibit zero covariance (Section 3.6). The $4\times 4$ case demonstrates that this property does not depend on the existence of a central cell at the origin. The algebraic content—that equal row and column sums imply vanishing covariance—is purely a statement about weighted sums and holds regardless of grid parity.
@@ 5.9 Algorithmic Implications
The energy landscape perspective suggests natural algorithmic strategies for magic square generation beyond classical backtracking. Since magic squares are isolated local minima, gradient-based descent in the $E(S)$ objective becomes viable. The challenge lies not in navigating a rugged landscape (entropy is modest and roughly constant at $H\approx 4.4$), but in the sheer breadth of the search space and the rarity of solutions.
@@ 6 Conclusion
Our central theoretical contributions are the Zero-Covariance Theorem and its energy landscape interpretation. The theorem establishes that an arrangement of integers in a square grid is a magic square if and only if the covariance between spatial position and assigned value vanishes. The energy characterization goes further: magic squares are precisely the ground states of a natural energy functional measuring geometric imbalance. These characterizations transform the discrete, combinatorial magic property into continuous conditions with physical meaning.
For orders $n=3$ and $n=4$, we have provided complete proofs of the Zero-Covariance Theorem, verified exhaustively for $n=3$ through enumeration of all 362,880 permutations. The Energy Characterization identifies magic squares as isolated local minima in the covariance landscape, with our perturbation analysis confirming that every single-swap perturbation strictly increases the energy. For general $n$, we state the result as a conjecture supported by extensive computational evidence.
@@
The Magic Gem framework opens several directions for future research. The geometric classification of magic squares, particularly for $n=4$ where 880 essentially different squares exist, could complement traditional algebraic approaches. Extensions to magic hypercubes, semi-magic squares, and pandiagonal squares would test the generality of our energy characterization. The scaling behavior of the perturbation gap—observed to decrease with nn—raises intriguing questions about the structure of large magic squares.
@@ A.1 Proof Strategy Overview
The Energy Characterization Theorem (Section 3.13) is established via two complementary approaches:
@@
Converse Direction (Exhaustive Verification): For $n=3$, exhaustive enumeration of all 362,880362{,}880 arrangements confirms that only magic squares achieve zero energy. For $n\geq 4$, large-scale sampling provides strong computational evidence.
@@ A.2 Detailed Proof: Magic Implies Zero Covariance
We provide a complete proof of Section 3.6, showing that every magic square has vanishing row and column covariances.
Detailed Proof of Section 3.6.
Let $S=(s_{ij})$ be an $n\times n$ magic square, and define coordinates as in Section 3.1:
@@ A.3 Why Row and Column Covariances Alone Are Insufficient
A.4 Energy Characterization Conjecture for General $n$
For $n\geq 4$, exhaustive enumeration becomes computationally infeasible ($16!\approx 2\times 10^{13}$ for $n=4$, and $25!\approx 10^{25}$ for $n=5$). We therefore state the general result as a conjecture, supported by extensive computational evidence.
Conjecture A.1 (Energy Characterization—General $n$).
For all $n\geq 3$, an arrangement $S$ of $\{1,2,\ldots,n^{2}\}$ in an $n\times n$ grid is a magic square if and only if the total magic energy vanishes:
$E(S)=\mathrm{Cov}(X,Z)^{2}+\mathrm{Cov}(Y,Z)^{2}+\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2}=0.$ (13)
This formulation is conceptually appealing: magic squares are precisely those arrangements that are simultaneously uncorrelated with all four structural directions of the grid (rows, columns, main diagonal, anti-diagonal).
Important clarification: Note that requiring all four covariances to vanish is essential. As discussed in Section 3.4.2, the conditions $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ alone do not characterize semi-magic squares—760 non-semi-magic arrangements satisfy these conditions for $n=3$. Only the complete energy functional $E(S)=0$ characterizes magic squares.
Supporting evidence:
  1. 1.
    Proven for $n=3$: Exhaustive enumeration of all 362,880362{,}880 permutations confirms that exactly 8 arrangements have $E(S)=0$, all of which are magic squares (the Lo Shu and its $D_{4}$ symmetry variants).
  2. 2.
    Strong evidence for $n=4$: All 7,040 known magic squares satisfy $E=0$ exactly. Large-scale random sampling ($6\times 10^{7}$ permutations) found zero non-magic arrangements with $E<10^{-10}$.
  3. 3.
    Strong evidence for $n=5$: All constructed magic squares satisfy $E=0$ exactly. Large-scale random sampling ($4\times 10^{8}$ permutations) found zero non-magic arrangements with $E<10^{-10}$.
  4. 4.
    Total verification: Over 460 million arrangements tested across $n=3,4,5$ with zero counterexamples observed.
Note that the sampling approach tests the forward direction robustly (confirming all magic squares have $E=0$) but can only provide probabilistic evidence for the converse (that all $E=0$ arrangements are magic). Given the extreme rarity of magic squares (probability $\sim 10^{-10}$ for $n=4$ and $\sim 10^{-17}$ for $n=5$), finding a counterexample by random sampling is unlikely even if one existed, making this evidence suggestive but not conclusive.
@@ A.6 Summary Statistics
Section 4 summarizes key statistics for magic squares of orders 3, 4, and 5.

### v2

@@ 5.1.1 The Vector Interpretation
This interpretation extends to the diagonal conditions. The diagonal covariances $\mathrm{Cov}(D_{\mathrm{main}},Z)$ and $\mathrm{Cov}(D_{\mathrm{anti}},Z)$ measure whether diagonal positions predict value. For magic squares, all four directional indicators—horizontal position, vertical position, main diagonal membership, and anti-diagonal membership—are uncorrelated with value. The low-mode energy $E_{\mathrm{low}}(S)$ is the sum of squared correlations with these four aggregate directions; the complete energy $E_{\mathrm{full}}(S)$ extends this to individual row and column indicators.
@@ 5.1.2 Physical and Algebraic Interpretations
The variances themselves admit exact closed forms: since the $z$-values are a permutation of $\{1-\mu,2-\mu,\ldots,n^{2}-\mu\}$ where $\mu=(n^{2}+1)/2$, a standard calculation yields $\mathrm{Var}(Z)=(n^{4}-1)/12$ and $\mathrm{Var}(X)=\mathrm{Var}(Y)=(n^{2}-1)/12$. Combined with the zero off-diagonal covariances, this completely determines the covariance matrix of the Magic Gem coordinates for any order $n$.
From an algebraic perspective, the zero-covariance condition is equivalent to the defining property of magic squares—equal row and column sums—expressed in the language of moments. The first moment (mean) of each coordinate vanishes by construction; the zero-covariance condition asserts that certain second moments (cross-terms) also vanish. In fact, this extends to higher-order moments: since $X$ depends only on column index $j$, we have $\mathbb{E}[X^{k}Z]=\frac{1}{n^{2}}\sum_{j}X_{j}^{k}\cdot(\text{column}_{j}\text{ sum}-n\mu)=0$ for any power $k$, because all column sums equal $M(n)=n\mu$. Similarly $\mathbb{E}[Y^{k}Z]=0$ for all $k$ follows from equal row sums. Thus magic squares exhibit vanishing cross-moments $\mathbb{E}[X^{k}Z]=\mathbb{E}[Y^{k}Z]=0$ at all orders—a stronger form of statistical balance than the covariance condition alone. This suggests a hierarchy: might there be “super-magic” squares characterized by additional vanishing moments beyond those implied by equal line sums?
@@ 5.2 Magic Squares as Ground States
The energy landscape perspective developed in Section 3.7 offers perhaps the most illuminating interpretation of our results. The complete covariance energy $E_{\mathrm{full}}(S)$ provides a natural “imbalance” measure for arrangements, and magic squares emerge as the unique configurations where this imbalance vanishes completely.
@@
The dynamical systems interpretation is equally rich. Each magic square defines a basin of attraction in the discrete configuration space, though these basins have measure zero. The comprehensive perturbation analysis (Figure 2) confirms that there are no “flat directions” around magic configurations: across 116,388 tested perturbations of 923 magic squares, every single swap increases the energy, supporting the conclusion that magic squares are strict local minima. Notably, while magic squares are the unique global minima (with energy zero), preliminary analysis for $n=3$ reveals additional local minima where no single swap decreases energy. Exhaustive enumeration finds 16 such non-global local minima under $E_{\mathrm{full}}$, and 336 under the simpler $E_{\mathrm{low}}$. This ruggedness has algorithmic implications: gradient-based search methods may become trapped in local minima, and the choice of energy functional affects the landscape structure.
The perturbation analysis reveals additional structure within the magic square population. For $n=3$, all eight magic squares exhibit exactly the same minimum perturbation gap ($\Delta_{3}=0.0988$), a consequence of the $D_{4}$ symmetry preserving the embedding geometry. More surprisingly, the 880 $n=4$ magic squares cluster into three distinct bands of perturbation resistance, suggesting the existence of geometric subclasses with different stability properties. This trimodal structure appears to correlate with convex hull geometry, indicating a connection between polyhedral structure and energy landscape stability. This finding warrants further investigation across the full $n=4$ equivalence-class population. The minimum gap decreases with nn—from 0.099 for $n=3$ to 0.004 for $n=4$ to 0.002 for n=5n=5—quantifying how the energy wells become progressively shallower as order increases. The precise scaling law requires investigation at additional orders.
@@ 5.3 Scaling Laws and Landscape Evolution
Despite the dramatic growth in configuration space ($n!^{2}$ arrangements), the landscape entropy remains within a narrow band ($\approx 4.3$$5.0$), decreasing from $n=3$ to $n=4$ and then stabilizing. This invariance suggests that the relative difficulty of local optimization is independent of scale: the ruggedness of the landscape, measured by the distribution’s shape rather than its position, does not fundamentally change with $n$. This has practical implications for algorithm design: local search heuristics that work for small orders may transfer to larger ones, with the primary challenge being the absolute energy scale rather than landscape structure.
Low-Mode Energy Landscape.
Across over 460 million sampled arrangements—including exhaustive enumeration for $n=3$, 60 million samples for $n=4$, and 400 million for n=5n=5—no zero-energy configurations of the low-mode energy $E_{\mathrm{low}}$ were discovered by random sampling. However, this does not imply that magic squares are the only zeros of $E_{\mathrm{low}}$ for $n\geq 4$; as shown in Remark 3.10, non-magic zeros exist but are too rare for random sampling to detect. The low-mode and complete energies measure different aspects of balance, and only the complete energy $E_{\mathrm{full}}$ (Theorem 3.17) provides a rigorous characterization for all orders. The probability of discovering a magic square by random sampling decreases super-exponentially with $n$, from roughly $1$ in 45,00045{,}000 for $n=3$ to about $2\times 10^{-17}$ for $n=5$ (roughly one in $5\times 10^{16}$). This quantifies the “needle in a haystack” intuition with concrete numbers.
@@ 5.4 Even versus Odd Orders
Despite these differences, magic squares of all orders exhibit zero covariance (Proposition 3.6). The $4\times 4$ case demonstrates that this property does not depend on the existence of a central cell at the origin. The algebraic content—that equal row and column sums imply vanishing covariance—is purely a statement about weighted sums and holds regardless of grid parity.
@@ 5.9 Algorithmic Implications
The energy landscape perspective suggests natural algorithmic strategies for magic square generation beyond classical backtracking. Since magic squares are isolated local minima, gradient-based descent in the $E_{\mathrm{full}}(S)$ objective becomes viable. The challenge lies not in navigating a rugged landscape (entropy is modest, $H\approx 4.3$$5.0$), but in the sheer breadth of the search space and the rarity of solutions.
@@ 6 Conclusion
Our central theoretical contribution is the Complete Energy Characterization Theorem (Theorem 3.17), which establishes that an arrangement of integers in a square grid is a magic square if and only if its covariance with every row, column, and diagonal indicator vanishes. Equivalently, magic squares are precisely the ground states of the complete energy functional $E_{\mathrm{full}}$, valid for all orders $n\geq 3$. This characterization transforms the discrete, combinatorial magic property into a statistical orthogonality condition with physical meaning.
The proof is elementary once the right framework is established: each row/column covariance vanishes if and only if that line sums to $M(n)$, and similarly for the diagonals. The perturbation analysis confirms that magic squares are isolated local minima, with every single-swap perturbation strictly increasing the energy.
We also investigated a natural simplification—the four-term “low-mode” energy using aggregate position coordinates $X$, $Y$ and diagonal indicators. This characterizes magic squares for $n=3$ (verified exhaustively), but defines a strictly larger class of “low-mode balanced” arrangements for $n\geq 4$. The existence of non-magic zeros of the low-mode energy, while initially surprising, follows from a dimensional argument: four linear constraints cannot determine all $2n$ line-sum conditions. Explicit counterexamples (Remark 3.10) and comparative energy analysis (Figure 4) illuminate this distinction, highlighting the necessity of the complete row/column indicator basis for a rigorous characterization.
@@
The Magic Gem framework opens several directions for future research. The geometric classification of magic squares, particularly for $n=4$ where 880 essentially different squares exist, could complement traditional algebraic approaches. Extensions to magic hypercubes, semi-magic squares, and pandiagonal squares would test the generality of our framework. The “low-mode balanced” class identified by $E_{\mathrm{low}}=0$ for $n\geq 4$ is itself an interesting mathematical object worthy of further study.
@@ A.1 Proof Strategy Overview
The Complete Energy Characterization Theorem (Theorem 3.17) is established via two complementary approaches:
@@
Converse Direction: For the complete energy $E_{\mathrm{full}}$, the converse follows directly from the row/column covariance characterization (Proposition 3.15, Proposition 3.12). For the low-mode energy $E_{\mathrm{low}}$, the converse holds only for $n=3$ (verified by exhaustive enumeration of all 362,880362{,}880 arrangements).
@@ A.2 Detailed Proof: Magic Implies Zero Covariance
We provide a complete proof of Proposition 3.6, showing that every magic square has vanishing row and column covariances.
Detailed Proof of Proposition 3.6.
Let $S=(s_{ij})$ be an $n\times n$ magic square, and define coordinates as in Definition 3.1:
@@ A.3 Why Row and Column Covariances Alone Are Insufficient
A.4 Complete vs. Low-Mode Energy Characterization
The main text establishes two energy functionals with different characterization properties.
A.4.1 Complete Energy (Theorem)
The complete magic energy uses individual row and column indicator covariances:
$E_{\mathrm{full}}(S)=\sum_{k=0}^{n-2}\mathrm{Cov}(R_{k},Z)^{2}+\sum_{\ell=0}^{n-2}\mathrm{Cov}(C_{\ell},Z)^{2}+\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2},$ (17)
where $R_{k}(i,j)=\mathbf{1}\{i=k\}$ and $C_{\ell}(i,j)=\mathbf{1}\{j=\ell\}$.
Theorem A.1 (Complete Energy Characterization—All $n$).
For all $n\geq 3$, an arrangement $S$ of $\{1,2,\ldots,n^{2}\}$ in an $n\times n$ grid is a magic square if and only if $E_{\mathrm{full}}(S)=0$.
This is proven in Theorem 3.17 of the main text. The proof follows directly from the row/column covariance characterization: each row covariance vanishes iff that row sums to $M(n)$, each column covariance vanishes iff that column sums to $M(n)$, and each diagonal covariance vanishes iff that diagonal sums to $M(n)$.
A.4.2 Low-Mode Energy (Partial Characterization)
The simpler low-mode energy uses aggregate position coordinates:
$E_{\mathrm{low}}(S)=\mathrm{Cov}(X,Z)^{2}+\mathrm{Cov}(Y,Z)^{2}+\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2}.$ (18)
For $n=3$: Exhaustive enumeration confirms that $E_{\mathrm{low}}(S)=0$ if and only if $S$ is a magic square. Exactly 8 of 362,880 arrangements achieve zero energy, all being the Lo Shu and its $D_{4}$ symmetry variants.
For $n\geq 4$: The condition $E_{\mathrm{low}}(S)=0$ defines a strictly larger class of “low-mode balanced” arrangements that properly contains the magic squares. Counterexamples exist: for example, the $4\times 4$ arrangement
$\begin{pmatrix}9&4&12&3\\ 16&1&14&7\\ 10&15&11&8\\ 2&5&6&13\end{pmatrix}$
has $E_{\mathrm{low}}(S)=0$ (both diagonals sum to 34, and the first-moment conditions are satisfied), but row sums $(28,38,44,26)$ and column sums $(37,25,43,31)$ are not equal. This demonstrates that the four-term energy is insufficient for $n\geq 4$.
Why random sampling failed to detect counterexamples: The probability of a uniformly random permutation achieving $E_{\mathrm{low}}=0$ is extremely low (estimated below $10^{-8}$ for $n=4$), so 460 million samples are unlikely to encounter such configurations. However, targeted search methods (e.g., hill-climbing on the constraint residuals) readily find them. This illustrates a fundamental limitation of random sampling for disproving “if and only if” conjectures.
@@ A.6 Summary Statistics
Table 4 summarizes key statistics for magic squares of orders 3, 4, and 5.